OpenStudy (anonymous):

How big is a ton? That is, what is the volume of something that weighs a ton? To be specific, estimate the diameter of a 1-ton rock, but first make a wild guess>: will it be 1 ft across, 3 feet, or the size of a car? (Hint: Rock has mass er volume about three times that of water, which is 1 kg per liter(10^3cm^3) or 62lb per cubic foot)

OpenStudy (anonymous):

This is a terrible question. A ton is a measure of mass, not volume or density. A short ton is 2000 lb. A metric tonne is 1000 kilograms. Taking the rock to be spherical, the volume is expressed as \[{4 \over 3} \pi r^3\]With the given density value, we can say that\[m = \rho \cdot {4 \over 3} \pi \cdot r^3\]where \(\rho\) is the density.

OpenStudy (unklerhaukus):

one cubic meter of water equals 1 tonne \[1m^3 \times H_2O\] 1 ton is about 1.1tonne